if (e^x+e^-x) / 2 = 1 then find e^x
Can you please me out? Thanks so much in advance:)

Answer by jim_thompson5910(35256)

(e^x+e^-x) / 2 = 1

(e^x+1/e^x) / 2 = 1

((e^x)^2+1)/(2e^x) = 1

Let z = e^x

((e^x)^2+1)/(2e^x) = 1

(z^2+1)/(2z) = 1

z^2+1 = 2z

z^2-2z+1 = 0

(z-1)^2 = 0

z-1 = 0

z = 1

e^x = 1

Final Answer: e^x = 1